### 19 Dec 2023 v1

# Designing Guiding Principles for NLP for Healthcare: A Case Study of Maternal Health

ABSTRACT
Objective: An ethical framework for the use of large language models (LLMs) is urgently needed to shape how natural language processing (NLP) tools are used for healthcare applications. Drawing directly from the voices of those most affected, we propose a set of guiding principles for the use of NLP in healthcare, with examples based on applications in maternal health.
Materials and Methods: We led an interactive session centered on an LLM-based chatbot demonstration during a full-day workshop with 39 participants, and additionally surveyed 30 healthcare workers and 30 birthing people about their values, needs, and perceptions of AI and LLMs. We conducted quantitative and qualitative analyses of the interactive discussions to consolidate our findings into a set of guiding principles.
Results: Using the case study of maternal health, we propose nine principles for ethical use of LLMs, grouped into three categories: (i) contextual significance, (ii) measurements, and (iii) who/what is valued. We describe rationales underlying these principles and provide practical advice.
Discussion: Healthcare faces existing challenges including the balance of power in clinician-patient relationships, systemic health disparities, historical injustices, and economic constraints. Our principles serve as a framework for surfacing key considerations when deploying LLMs in medicine, as well as providing a methodological pattern for other researchers to follow.
Conclusion: This set of principles can serve as a resource to practitioners working on maternal health and other healthcare fields to emphasize the importance of technical nuance, historical context, and inclusive design when developing LLMs for use in clinical settings.
1 OBJECTIVE
As generative natural language processing (NLP) methods and large language models (LLMs) have increased in size and performance, so has excitement increased about their clinical use [34]. However, due to the sensitivity of health-related data and the potential for harmful negative outcomes, employing NLP technologies in clinical settings requires careful consideration. Healthcare involves numerous stakeholders who can provide useful contributions towards the design of these technologies, e.g., care-seekers, clinicians, researchers, and others all have a shared interest in the development and use of these technologies.
In this work, we identify a set of guiding principles that support well-grounded and ethical uses of NLP technologies in healthcare, developed with input from diverse stakeholder groups. We employ a participatory design framework [27] to amplify the voices of stakeholder groups who have previously been underrepresented in the discussion around use of NLP technologies in healthcare.
We develop our guidelines by focusing on a case study of maternal healthcare in the United States. We choose maternal health for three reasons. First, pregnancy and childbirth are common events that often comprise a person’s sole or major interaction with the healthcare system. Second, there are many prior research studies and applications of NLP methods focused on maternal healthcare [23, 17, 21, 35]. Third, maternal health is a “perfect storm” of healthcare vulnerabilities, with historical biases and power dynamics influencing care [12, 25]. Studying maternal health allows us to map our investigation and recommendations over a set of specific and grounded opportunities and risks.
We solicited key considerations about an LLM-based chatbot and NLP tools from diverse stakeholders, including clinicians, birthing people, researchers, community health workers, government and non-profit workers, and others. We collected perspectives in two main ways (Figure 1). First, we introduced stakeholders to NLP technologies, allowing them to surface concerns, opportunities, and discussion through interaction with these technologies in a controlled setting. Second, we conducted surveys to gather information about questions, priorities, and risks.
Through analysis of the collected data, iterative rounds of discussion, and literature review, we consolidated and identified nine guiding principles for the use of NLP in healthcare, organized according to three themes: context, measurements, and values. These themes are inspired by our maternal health case study but provide a transferable blueprint for mapping tensions in other healthcare areas, supporting NLP practitioners and healthcare administrators on decisions related to the design and deployment of NLP tools.
2 BACKGROUND AND SIGNIFICANCE
Ethical guidelines for NLP/AI and healthcare
Prior work has developed guidelines and frameworks for the use of artificial intelligence (AI) for healthcare, and provide important guidance for AI practitioners. These works take a general view of healthcare and are often focused on ML practitioners and their processes. For example, Chen et al. [9] and Wiens et al. [37] provide an overview of the NLP development pipeline and recommendations focused on each pipeline step. The recommendations in these works are based on literature reviews and healthcare examples.
We build on these works by focusing our investigation on a specific healthcare context and direct feedback from its stakeholders, then drawing broader lessons for the rest of healthcare. In particular, we elicit values from various stakeholders using a framework introduced by Jakesch et al. [20], which aims to model the “human process” at the center of healthcare [16]. As recommended by Rajkomar et al. [30] and following prior work [36], we follow research process that draws on participatory design, a collaborative design approach that involves stakeholders and end users in the design process for new technologies [27], such that their needs drive design decisions and result in tools that better address their concerns and challenges [14, 22, 31, 19]. We also make use of participatory design in the formation of our guiding principles; uniquely, we survey stakeholders connected to a specific healthcare topic to better illuminate tensions and priorities for NLP usage for healthcare.
Maternal health: vulnerabilities and NLP applications
Rates of pregnancy-related deaths and complications have increased over the last thirty years in the U.S., with particularly high rates among Hispanic and Black birthing people and non-U.S. citizens [5, 11, 12, 24, 25]. For example, Black women are three times more likely to die in childbirth than white women [12, 25]. Most frustratingly, 46% of maternal deaths of Black women and 33% of maternal deaths of white women are estimated as preventable [6], but measurements of these problems are challenging [13].
Such urgent problems need every tool in the toolbox, and NLP models can facilitate the analysis of large-scale datasets. Prior work has used NLP methods to study the prevalence of adverse outcomes via social media data [23], extract lactation information from drug labels [17], predict high-risk pregnancies from electronic health records (EHR) [21], examine how news media discuss pregnancy and exposure to isotretinoin [26], and assist in analysis of phone interviews about breastfeeding [35]. In our review of this prior work in NLP, we observe the following patterns: (i) many studies attempt to surveil and predict psychological states of the birthing person, e.g., predicting postpartum depression, and (ii) most studies use either social media or EHR data, with a bias in studies published at NLP venues towards social media data produced by birthing people.
3 MATERIALS AND METHODS
In Figure 1, we provide an overview of our methods, which center around a structured interface for participants to query an LLM-based chatbot. After participants have interacted with this demonstration, they move to virtual breakout discussion sessions led by volunteer moderators, and a survey, in which we asked participants to express their perceptions of NLP tools for maternal healthcare. Non-workshop participants wrote individual responses to our discussion questions rather than worked together with a moderator. Our study was approved by the IRB at the Allen Institute for AI.
Figure 1: Study overview, including the three participant cohorts, the chatbot demonstration, and the surveys and discussions followed by analysis and design of the guiding principles.
3.1 Participant Cohorts
We ran this study with three participant cohorts. Using the Prolific [^1] https://www.prolific.com/ survey platform, we recruited one cohort of birthing people (gave birth in the last five years) and one cohort of healthcare workers (have ever worked in any healthcare profession). We required that all participants be over 18 years old and be located in the U.S.
For the third cohort, we led a live, virtual session with a group of participants who received training about LLMs and their risks. Our session was part of a virtual workshop hosted by the Association of American Medical Colleges (AAMC) on NLP and maternal health in which participants were introduced to the basics of NLP, applications of NLP to maternal health, and biases in NLP tools. Most but not all of these participants also worked in healthcare, though compared to our cohort from Prolific, these participants more often worked in community and research roles rather than as clinicians.
For full descriptions of these cohorts, see Table 1 in §4.1.
3.2 Design of chatbot demonstration
We built an interface wrapper (Figure 2) around GPT-3.5 from OpenAI. [^2] Compared to GPT-4, which was also accessible at the time of our study, GPT-3.5 is less expensive and has lower rate limits, which was especially important for the workshop setting, where 100+ people were interacting with the bot at any moment. Our goal was to demonstrate high performance for the bot rather than create a reproducible system; otherwise, an open model would have been preferable. Our goal with this demonstration was to give participants a guided but unconstrained interaction with a chatbot, gathering their queries, structured feedback, and perceptions related to the chatbot and maternal health. For many of our participants (see Figure 6 in the Appendix), this was their first direct interaction with a modern LLM. We included prominent warnings about (i) hallucinations and unreliability, (ii) data collection (see Figure 2), and (iii) the importance of always seeking the advice of a healthcare professional. We also provided examples of the kinds of maternal health questions one might ask a chatbot.
Figure 2: Screenshots of our chatbot demo. Responses were generated using GPT-3.5 from OpenAI. Users could ask multiple questions about maternal health. The workshop demonstration additionally included a free-text entry box for participants to add descriptions of situations in which they might use the chatbot.
3.3 Design of discussion sessions
Workshop participants
Workshop participants were divided into eight breakout rooms and participated in a 30-minute free-form discussion session, moderated by workshop organizers and trained volunteers. Discussion sessions were recorded after obtaining consent from all participants. During the session, participants were shown the following discussion prompts:
  • How was your experience with the chatbot? What stood out to you about the responses?
  • What are your dream NLP tools for maternal health? What tools should never be built?
  • Which maternal health stakeholders (birthing people, nurses, doulas, etc.) would benefit or be hurt by NLP tools?
  • What principles should guide the use of NLP for maternal health? What should be the goals and guardrails?
We received video and audio recordings for four out of eight sessions as well as written comments from five sessions, which we included in our analysis.
Healthcare workers and birthing people
Non-workshop participants first interacted with our chatbot demonstration, then wrote answers to our discussion questions as part of a followup survey (like the workshop participants).
3.4 Survey design
Our goals in the survey were to (i) elicit participants’ perceptions of NLP tools, (ii) learn about participants’ information seeking goals, and (iii) build collaborative rankings of values that should guide NLP uses for maternal health. Surveys given to each cohort were identical except that for birthing people, questions about career were replaced with questions about whether the participant would like their healthcare team to use and/or disclose their use of AI tools. Figure 3 provides a condensed view of the survey we gave to participants; see Appendix C for the full set of survey questions. Prolific workers were paid an average of \$15/hour.
Figure 3: Simplified view of the survey given to participants. The question about trust is drawn from Baughan et al. [3], while the questions about general AI attitudes are drawn from a frequently reused set by Nickell and Pinto [28]. The listed values were taken from Jakesch et al. [20]; definitions were provided for each value. Questions in the final box were shown only to the cohort of birthing people.
4 RESULTS
Cohort Race/Ethnicity Age Highest Education Gender
Workshop Participants ($N=39$)
38% community nonprofits
27% pop./public health research
24% comm. health/promotara
24% local/state public health
19% HC management/admin
16% HC services researcher
11% other peri. HC provider
11% other non-HC peri. support
8% doula
8% non-peri. HC provider
13.5% all other groups
41% Afr-Amer/Black
41% White
16% Hispanic/Latino/a/x
5% South Asian
19% all other groups
35% 35-44
30% 25-34
19% 45-54
11% 55-64
5% 65-74
38% MS, MPH, etc.
30% PhD
24% BA, BS, etc.
11% all other groups
92% women
5% men
3% no answer
0% non-binary
Healthcare Workers ($N=30$)
20% nurse
17% pharmacy
10% physician
10% medical tech
10% medical assistant/aide
10% research
23% all other groups
33% have worked in maternal/perinatal healthcare
57% White
23% Afr-Amer/Black
7% East Asian
7% Southeast Asian
9% all other groups
33% 35-44
30% 25-34
10% 18-24
3% 65-74
3% 55-64
50% BA, BS, etc.
17% MS, MPH, etc.
17% Trade School
10% MD, DO, etc.
7% Community College
6% all other groups
79% women
21% men
0% non-binary
Birthing People ($N=30$)
20% have worked in healthcare
7% have worked in maternal/perinatal healthcare
73% White
20% Hispanic/Latino/a/x
17% Afr-Amer/Black
12% all other groups
53% 25-34
37% 35-44
10% 65-74
33% BA, BS, etc.
30% High school or GED
13% Community College
10% MS, MPH, etc.
10% Trade School
7% PhD
3% Prof. Degree
97% women
7% non-binary
0% men
Table 1: Demographic description of participant cohorts. Healthcare workers and birthing people were recruited from the Prolific platform while the workshop participants took part in multiple training and educational sessions to learn about LLMs, their applications to maternal health, and their risks. Work category, race/ethnicity, highest education, and gender all allowed multiple selections, so the percents for these categories might not sum to 100%.
4.1 Survey Results
We release our survey results publicly to support future research. [^3] We remove demographic features for the workshop participants, as these participants interacted with one another and could potentially identify others’ survey responses. Participants consented to the sharing of their responses. All survey data is available in supplementary materials and at this public Github repository: https://github.com/maria-antoniak/maternal-health-principles Table 1 provides a summary of participant demographics and background. The workshop participants tended to represent the non-profit, community health, research, and governmental sectors in contrast to clinicians such as physicians and nurses that were more common in the Prolific cohort of healthcare workers. The workshop participants represented a more even distribution across racial/ethnic groups, but all cohorts had little representation in the East Asian, South Asian, and Southeast Asian groups. Overall, workshop participants reported less experience with AI and NLP than birthing people and healthcare workers recruited from Prolific (see Appendix A).
Figure 4 shows how frequently each value was selected by each cohort. Overall, safety, privacy, and performance were selected more often by birthing people and healthcare workers in comparison to workshop participants, who were more likely to select inclusiveness. The birthing people and workshop participants were united in being more likely to select human autonomy than the healthcare workers, and less likely than the healthcare workers to select transparency and accountability. Figure 5 compares coocurrences of values selected by different cohorts. Birthing people were more consistent in their selections, while healthcare workers had more variation in their selections. Overall, we do not find large differences in the value selections across stratifications for generalized trust, trust in healthcare workers, or generalized AI perceptions (see Appendix A).
Figure 4: Frequencies of selected values. Each participant was asked to select five values from a list of 12 curated by Jakesch et al. [20]. Participants were given definitions of the values, also drawn from Jakesch et al. [20]. Birthing people and healthcare workers overall responded more similarly to each other than to the workshop participants, but this was not always the case (e.g., transparency, human autonomy).
Figure 5: Co-occurrences of selected values. Each participant was asked to select five values from a list of 12 curated by Jakesch et al. [20]. Participants were given definitions of the values, also drawn from Jakesch et al. [20]. Overall, the workshop participants were less consistent in their selection of values than the birthing people and healthcare workers.
4.2 Thematic analysis of discussion sessions
We identified themes and concepts frequently brought up by study participants in workshop breakout discussions, written comments, free-text survey answers, and chatbot queries. We first performed open coding on responses from each source to identify relevant concepts, then axial coding to group these concepts under broad themes [10]. The authors collectively discussed these themes and iterated several times before synthesizing them into the final set of nine guiding principles presented in §4.3. We summarize these principles in the following section and include anonymized quotes from our study participants to highlight stakeholder pain points and further support the inclusion of each principle. Additionally, we provide actionable recommendations to NLP researchers and practitioners based on our own experience and expertise as NLP researchers and healthcare workers (Appendix B). Our thematic analyses also raised several additional points, which we discuss in §5.
4.3 Guiding Principles
Drawing from the discussion themes, our survey results, and related work (both applications work and high-level ethics guidelines like those by Chen et al. [9] and D’Ignazio and Klein [15]), we develop guiding principles for the responsible use of NLP in maternal health. Detailed descriptions and actionable recommendations derived from each principle are provided in Appendix B.
We present these principles grouped under three themes: context, measurements, and values. Context principles ask stakeholders to incorporate the fundamentals and history of maternal health in their applications, measurement principles discuss what to optimize for and how to evaluate, and value principles address how NLP researchers and practitioners should situate user voices and data relevant to our systems. We summarize each principle below and highlight in participants’ own words their importance. In quotes, workshop participants are attributed as W1-10, birthing people as B1-11, and healthcare workers as H1-15.
Theme 1: Context
Be aware of power dynamics in the care team  Maternal healthcare has a long and fraught history of shifting power dynamics in the care team [4]. These historical shifts, like male OB/GYNs replacing midwives in the early 20th century, can lead to both good and bad effects for the birthing person. NLP practitioners should know this history and work to improve rather than exacerbate existing hierarchies, being mindful of the impact of tools and automation on midwives, doulas, and other care workers whose placement is already precarious. “AI should not be used instead of actual and necessary medical care” (B1). “If mothers relied too heavily on AI instead of seeking professional help then the nurses and doulas may see fewer people seeking care” (B11).
Know the politics and implications of your measurements  Using NLP tools for data collection can have unintended consequences, e.g., (i) overemphasis on data collection can draw resources away from building solutions, (ii) reproductive health is politicized, and measurements may be used as evidence in support of unexpected agendas, and (iii) focusing only on measuring problems can contribute to “deficit narratives” that blame communities for their own challenges. “It’s key that the AI is actually able to provide solutions to problems that can be fixed, and not just simply acknowledging them” (H4). We therefore recommend clarifying the impact goal of an NLP project before focusing on what data to collect. Financial motivations also muddy these decisions. As one participant puts it, “I get worried about what’ll happen when insurance companies think there’s cost savings to using these tools…they can cut corners, have more profit…given the incentives in … the healthcare system” (W1).
Learn from healthcare traditions and communities  Thousands of years of cultural traditions, healthcare practices, and tool development already exist, and more recently, grassroots communities on the internet have formed to help birthing people prepare for and work through their challenges and experiences. “People have the mindset that this needs to actually replace things or that this needs to be the way in the future of addressing maternal health…diminishing all the work that’s been done in the past and the positives of having people that truly understand this work” (W2). Learning from traditions and communities can help avoid repeating mistakes made by others. “It [(AI)] should follow the same guidelines that medical professionals do: ‘Do no harm’ ” (H6).
Theme 2: Measurements
Optimize for outcomes that support the whole person  “It’s not just about the outcome, right? It’s about the whole experience” (W3). NLP research, like most AI research, is centered on structured tasks with defined and accessible outcomes, like maternal mortality rates. Reducing maternal mortality is important, but so are other outcomes, like achieving a positive birthing experience or breastfeeding without stress. As one participant says, “it would be great if AI could tell everything from conception to birth” (B10). Taking a more expansive view [^4] By “expansive view,” we do not mean that a single model must be capable of all tasks (as this may in fact be a weakness of current LLMs whose generalizability can lead to a lack of clear evaluation). Rather, we invite NLP researchers and practitioners to be open-minded about tasks they can address with their NLP work. of the tasks and optimizations that could benefit maternal health stakeholders can open the door to other opportunities. “A dream AI tool should have the ability to address and develop solutions for mothers who are struggling to cope with the demands of pregnancy and overall maternal health” (H4).
Fight against inequity and protect all groups  The healthcare system treats people differently on the basis of race, class, gender, etc. NLP models can mirror biases found in training datasets—and even worse, they can over-represent and exacerbate the impacts of those biases. “Everything that has the potential to benefit also has the potential to hurt” (W4). When evaluating NLP systems, one should consider both disparities in impact (health outcomes for different groups) and disparities in treatment (whether people with the same complaint receive different treatment). Tools should “allow providers to improve their own cultural competency or race-consciousness” (W5), and empower “those lacking social and financial support,” such as young mothers (H14).
Hold onto the human: empathy, emotion, relationships, complexity  NLP tools should account for human elements like emotion, empathy, and storytelling [8], remembering that each birth is unique and each birthing person has a unique set of circumstances, experiences, and preferences. “Human connection is important and should be emphasized so that better quality service can be provided” (W6); “your own judgment and/or human compassion components, wisdom, experience [are a] part of the care” (W7). Don’t shy away from outliers and difficult, subjective tasks. Sometimes the goal is not to solve a problem, but to better understand it. “Birth is so complicated though that I don’t see women ever replacing medical care with AI advice” (B11).
Theme 3: Values
Include the voices of those seeking care  The voices of the most vulnerable stakeholder group, birthing people, should be included in the design and development of NLP tools. While it may not always be possible to include birthing people on a research or production team, it is important to integrate their perspective throughout design and deployment processes through surveys and user studies. As one participant puts it, “there needs to be community input, there needs to be representation in creating these tools” (W8).
Always center the agency and autonomy of the person seeking care  Maternal healthcare has an unfortunate history of abuse and disregard for the birthing person’s agency. Rather than working from a perspective of skepticism about the birthing person’s ability to make decisions for themselves, practitioners should respect the person’s agency over their own experience. How NLP tools are being used to make decisions in maternal healthcare should be disclosed to the birthing person, such that “the person using it knows what kind of information or advice it can and cannot give” (B2). Correspondingly, some tools should not be built, such as anything “infringing on the rights of the mom or baby” (B3).
Respect and support your data sources  Generations of birthing people have passed down oral stories, written books, and created online content about pregnancy, labor, and the postpartum period. NLP studies and tools benefit from this knowledge and data, and should give credit and be designed to avoid supplanting systems of support that are already thriving. As several workshop participants observed, ChatGPT and other similar chatbots “don’t really note their references of where they’re getting the information from” (W9). “The principle that should guide these tools is to… have transparency for its sourcing of data” (H2).
5 DISCUSSION
We asked over 100 people from diverse groups about their values, practices, and perceptions of AI and chatbots. Using maternal health as a test case, we uncovered a set of guiding principles with lessons for all uses of NLP for healthcare. Our discovered themes touch on questions of power and agency, trust between health specialists and birthing people, and contextualizing innovations within a history of disenfranchisement and mistakes. Our study also opened further questions, which we discuss below.
Applicability to other healthcare settings
Maternal health served us as a representative health case study through which we could apply a focused lens to the use of LLMs for healthcare. Related healthcare settings include resource shortages in primary care [7], individuals seeking health information to diagnose symptoms [2], and racial considerations in disease severity models [29]. In each case, our guiding principles can help stakeholders think through whether and how to incorporate NLP tools into healthcare settings.
Weighing the risks and benefits of using NLP and LLMs in healthcare settings
Participants reported largely positive attitudes towards NLP tools with important caveats. Workshop participants unanimously agreed that NLP tools for maternal health should always be designed to be assistive rather than autonomous (holding on to the human) while having high hopes about the development and adoption of NLP tools in healthcare.
Settings in which NLP tools were viewed to be particularly useful to healthcare providers included: (i) reducing administrative burden, including better/faster communication and coordination of care across medical departments, and (ii) improving medical education and training, including assisting staff in developing better cultural sensitivity. On the other hand, tools for decision-making settings were viewed negatively. These results contrast with current trends in NLP research (§2), which tend towards surveillance of birthing people and tasks like risk prediction.
Birthing people viewed NLP tools very positively, especially for providing information and recommendations (centering their agency and autonomy). For example, participants wished they had the chatbot during their pregnancy: “I wish it [(AI)] had been around when my son was a newborn so I could interact with it during late night feedings. One, to give me something to do, and two, to make me feel like I wasn’t alone” (B12). In particular, birthing people emphasized how comforting it would be to have a fast and convenient resource to assuage fears about whether what they are experiencing is normal: “Just to have something there to ask questions to when I am not sure as to what is happening or when I need a quick answer” (B9). However, some participants cautioned that such tools should provide additional context to help patients understand whether the suggestions or advice are applicable to their specific situation due to differing levels of medical understanding.
Improving the utility of current LLMs
While study participants were impressed the chatbot’s responses and capabilities, they identified several areas for improvement, in terms of both response presentation and response content. In addition to wanting more attribution of data sources (respecting data sources), participants thought response presentation could be improved by (i) reducing constant use of “safeguards,” which hampers readability (e.g., “I am an LLM…”, “please consult a doctor…”, etc.) and (ii) tailoring responses to match a user’s literacy and reading levels, particularly when presenting complex medical recommendations. This further motivates recent work on generating plain language summaries [1] and personalized rewrites of medical information [18], and points away from a general-purpose chatbot. Discussants also suggested (i) incorporating cultural nuance in model responses (e.g., personalizing responses for the user’s cultural context and away from a single dominant culture [32, 33]) and (ii) improving model sensitivity to minor changes in wording, which often generated very different responses.
LLMs as part of a larger ecosystem to meet information needs
Around half of the birthing people and two healthcare workers compared their experience with the chatbot to their past experiences with internet search engines. Most framed the comparison favorably, writing “Often times people will google questions and try to sift through all the search results to find the applicable information. AI could make that a much more efficient process” (H9) and “It would be nice to be able to type in worries and fears to an AI bot and get accurate answers instead of going down rabbit trails on search engines that leave you more concerned” (B11). On the other hand, one participant noted dangers to LLMs similar to known dangers of popular healthcare websites: “People already diagnose themselves on WebMD. Providing more tools can be dangerous” (B4). These participants viewed LLMs as one more tool in an existing information ecosystem, carrying many of the same risks and benefits.
Limitations of our study protocol
The birthing people in our study were from the U.S., spoke English, and lack representation among Asian and Pacific Islander groups. We used ChatGPT-3.5 for the demonstration, since its rate limits fit the constraints of the live workshop, though the model is regularly updated and results may not be replicable.
Differences in results between our participant cohorts are likely driven by at least three factors, though we cannot make any conclusive causal claims: (i) the groups represent different stakeholders; (ii) the workshop participants were primed by their multi-hour exposure to lectures about NLP and AI biases; and (iii) the workshop participants opted into the workshop, which was themed around “Maternal Health Equity,” while the Prolific participants were solicited online and paid for their work. This shows the importance of engaging with multiple groups in different settings and integrating their viewpoints, not by enforcing agreement but by surfacing the inherent tensions and different viewpoints that exist in the perceptions of these technologies.
6 CONCLUSION
In consultation with healthcare professionals, birthing people, and other stakeholders, we developed a set of guiding principles for the use of NLP and LLMs in healthcare. This work serves as a guide to researchers and practitioners broadly on how to engage with affected peoples in healthcare contexts. We hope that researchers and clinicians will work with the people most affected—the people seeking care—and will read and follow the principles derived by our methods. Machine learning researchers should also explore how their tools can actively broadcast these principles or how model development pipelines can be adapted to better meet the needs of all stakeholders.
7 ACKNOWLEDGEMENTS
The authors would like to thank the many study participants for their engagement and responses and the Association of American Medical Colleges (AAMC) volunteers who moderated the live workshop sessions. We would also like to thank Qian Yang, Tal August, and Katherine Antoniak for their research input and feedback, as well as Arnavi Chheda, Michael Schmitz, Johann Dahm, and Michael Wilson for their technical support. This work was supported by the AAMC Center for Health Justice in collaboration with the Allen Institute for AI (AI2).
Appendix A Extended Survey Results
Figure 6: Answers to the survey question [Before this workshop,] How familiar were/are you with NLP, machine learning, and/or AI? Overall, the workshop participants were less familiar with these topics than the healthcare workers and birthing people who were recruited via the Prolific platform.
Figure 7: Answers to the four survey questions about general perceptions of AI. Overall, the workshop participants less frequently reported positive perceptions than the healthcare workers and birthing people who were recruited via the Prolific platform.
Figure 8: Answers to the two survey questions about trust. Overall, the healthcare workers were more trusting of healthcare providers.
Figure 9: Answers to the survey questions about who would benefit from an AI/NLP chatbot in their work, with the 5-point Likert scale score shown on the x-axis.
Figure 10: Selection of values broken apart by cohort and general trust. Positive indicates scores $>3$.
Figure 11: Selection of values broken apart by cohort and trust in healthcare providers. Positive indicates scores $>3$.
Figure 12: Selection of values broken apart by cohort and general perceptions of AI. Positive indicates scores $>3$.
Appendix B Extended Guiding Principles
Extended definitions of guiding principles and application examples are shown in Table 2 (Context), Table 3 (Measurements), and Table 4 (Values).
Description
Applications
Principle 1: Be aware of power dynamics in the care team.
Maternal healthcare has a long and fraught history of shifting power dynamics in the care team. These historical shifts, like male OB/GYNs replacing midwives in the early 20th century, can lead to both good and bad effects for the birthing person. While technology isn’t a solution to human power imbalances, it’s possible that the introduction of NLP tools could further influence these dynamics, and so NLP practitioners should know this history and work to improve rather than exacerbate existing hierarchies. In particular, NLP practitioners should be mindful of the impact of their tools on midwives, doulas, and other care workers whose placement is already precarious.
Understand historical power shifts in the care team. Consider carefully who you expect to use your tool and who will be impacted (or potentially replaced) by your tool. What kinds of unintended shifts could occur? Consider these questions: Whose data and experiences are represented in the training of the tool? Who will provide predictions and advice? Who will receive predictions and advice? Who will have access to and control over the NLP tool? Are you concentrating more power within a single care team role? Reflect on your own position in the power hierarchy as an NLP practitioner collaborating with healthcare workers and birthing people. Who is making decisions about the design and application of NLP methods to maternal health?
Principle 2: Know the politics and implications of your measurements.
Reliable data can spur action, inspire the allocation of resources, and improve outcomes, but much maternal healthcare data is still missing. NLP tools can help fill these gaps by extracting information from texts or predicting missing data.
However, using NLP tools for data collection can have unintended consequences. (1) An overemphasis on data collection can draw resources away from building solutions. (2) Reproductive health is politicized, and your measurements may be used as evidence in support of unexpected agendas. (3) Focusing only on measuring problems can contribute to “deficit narratives” that blame communities for their own challenges.
Clarify your research and/or impact goals. Consider the current allocation of resources and where help is most needed when choosing problems to tackle. While additional data collection can be very useful, even for well-known problems, it can also draw resources away from other important problems. Think through possible narratives and take care about what data is collected, what measurements are produced, and how those measurements are communicated to the public. Know that your measurements and predictions could contain information that could be used to incriminate the birthing person in some places and situations. Similarly, this data could sometimes be used for unintended purposes, like personalized advertising and the setting of insurance rates.
Principle 3: Learn from healthcare health traditions and communities.
When designing NLP studies and tools, you don’t need to start from scratch. Thousands of years of cultural traditions, healthcare practices, and tool development already exist, and more recently, grassroots communities on the internet have formed to help birthing people prepare for and work through their challenges and experiences. Learning from these traditions and communities can help you avoid repeating mistakes made by others.
NLP is one tool in a big toolbox. An interdisciplinary team can help you avoid reinventing the wheel. Storytelling is an important way for communities and generations to collectively teach and learn about maternal health, as well as a way for the birthing person to process their experience. Consider how NLP tools can support and learn from storytelling rather than replacing this practice. Misinformation also exists in these datasets. Consider carefully how to balance handling misinformation—recognizing that the root cause is often mistrust in the healthcare system and your NLP study could contribute to that mistrust—while also valuing the expressions of personal experiences.
Table 2: Guiding principles about contextual significance.
Description
Applications
Principle 4: Optimize for outcomes that support the whole person.
Much existing NLP research for maternal health has focused on maternal mortality and postpartum depression but not on the many other challenges facing birthing people and their care teams. Maternal health also encompasses sexual health, family planning, pregnancy, the “fourth trimester” of postpartum health and recovery, breastfeeding, and more.
How we define success is important. NLP research, like most AI research, is centered on structured tasks with defined and accessible outcomes, like maternal mortality rates. Reducing maternal mortality rates is important, but so are other outcomes, like achieving a positive birthing experience or breastfeeding without stress.
Create new datasets with labels prioritizing other parts of maternal health. Consider who and what is not currently represented in existing datasets. Use social media and birthing person interviews to get a holistic overview of what outcomes birthing people care about. Run more long-term user studies to support a broader range of outcomes of interest. The outcome doesn’t have to be a number; be open to more subjective measurements. Adoption of your tool is also an outcome, and you don’t always have to move the needle on a traditional NLP metric.
Principle 5: Fight against inequity and protect all groups of people seeking care.
The healthcare system often treats birthing people differently on the basis of their race, class, gender, etc. These biases can spread into your collected datasets, influence your design process, and flow downstream into the impacts of your NLP study or tool.
While these issues exist throughout and beyond healthcare, maternal healthcare is particularly vulnerable to disparate treatment, and is an important place to challenge those disparities. (1) For many people, childbirth is their defining, most invasive, and riskiest experience with the healthcare system. (2) Numerous decision points require the birthing person and care team to work together to identify a pathway unique to their circumstances. (3) Some of these decisions rely on subjective evaluations of the birthing person’s experiences by the care team (e.g., feeling pain, bonding with the baby).
These qualities open the doorway to disparate treatment and heighten the long term impact of mistreatment.
When modeling outcomes, consider both disparate impacts and disparate treatments. Sometimes people with the same complaint aren’t given the same treatment (e.g., pain management disparities). But in other cases, applying the same treatment across the board, without regard for individual circumstances, is also an inequity. Remember that NLP models can mirror biases found in their training datasets—and even worse, they can overrepresent those biases, exacerbating the impacts of existing biases. When building tools that communicate with birthing people (e.g., chatbots), consider a diversity of vocabularies, accents, and languages. Beware counterfactuals as a methodological bandaid. Measuring and addressing disparities are important, but simply switching mentions of race or gender (e.g., replacing all mentions of she with he) ignores how living in society with those identities affects many aspects of one’s life and can lead to correlated differences in your data. Actively recruit birthing people from diverse groups to participate in your NLP study and/or give feedback on your NLP tool.
Principle 6: Hold onto the human: empathy, emotion, relationships, complexity.
Maternal healthcare revolves around meaning-laden events that can forge, strengthen, and test the human relationships between the birthing person, their child, their partner and family, their social network, and their healthcare team. We should build NLP tools that account for human elements like emotion and empathy, remembering that each birth is unique and each birthing person has a unique set of circumstances, experiences, and preferences. Individual stories and voices should be highlighted, not flattened; contradictions and outliers should be valued, not discarded.
Look for outliers and value what they can teach you, rather than removing them. A one-size-fits-all approach is probably not appropriate. Research and applications should prioritize personalization and/or allow the tool to be tailored by the individual birthing person or healthcare worker. Don’t be afraid to approach subjective tasks, like predicting satisfaction in the birthing experience, or tasks where inter-annotator agreement might be low, like labeling emotion-support in online birthing communities. Sometimes the goal isn’t to solve a problem but to better illuminate it. Include qualitative methods in your study design, complementing quantitative models with focus groups, interviews, grounded theory, etc. Support the expression of feelings and empathy by everyone involved. Optimize for giving the care team more space and time for empathy, conversation, and trust building.
Table 3: Guiding principles about measurements.
Description
Applications
Principle 7: Include the voices of those most effected.
Weight heavily the voices of people most affected by your tool. Center the principle of “nothing about us without us,” first popularized by disability activists. While the perspectives of NLP researchers and healthcare workers are also very important, the most vulnerable group — birthing people — must also be included in the design of your NLP tools and studies. The voices of birthing people, especially from marginalized groups, are often dismissed in healthcare, so aim to reduce barriers and distance between the birthing person and their care team.
Where possible, include birthing people and healthcare professionals in your research team. Include surveys and user studies throughout your research and design process Incorporate literature written by birthing people, and learn from related work about the birthing experience and birthing people’s perspectives and needs. This doesn’t mean putting all the responsibility on the birthing person or designing only for the birthing person. Part of supporting maternal health is encouraging support from the birthing person’s partner, family, and healthcare team. Be gender inclusive. Don’t automatically predict the gender of study subjects, and remember to include the concerns of trans people in your study design. Use gender inclusive language and follow the HCI Guidelines for Gender Equity and Inclusivity.
Principle 8: Always center the agency and autonomy of the person seeking care.
Maternal healthcare has an unfortunate history of abuse and disregard for the birthing person’s agency, and any research or tool-building for maternal healthcare must be conducted in awareness of this context. Just as healthcare workers should respect the bodily autonomy of birthing people, NLP systems should not be used to replace or circumvent the decision-making of birthing people. Any predictions should be transparently communicated, be reversible, and allow for recourse. Respect the birthing person’s experience; don’t work from a perspective of skepticism about the birthing person’s ability to make decisions for themselves. Birthing people are in charge of their experience.
Explore the construction of NLP tools that increase the agency available to the birthing person. For example, NLP tools might be used to provide birthing people with resources, explanations, descriptions of what to expect, or assistance for communicating with healthcare providers. If NLP tools are used to make or assist in clinical decisions, this should be disclosed to birthing people, and where possible, providers should obtain direct consent from birthing people. There should be a clear system for recourse if the birthing person is unhappy or confused about the use of an NLP system during their birthing process. Maintain privacy. Collect only necessary data and store the data securely. Avoid perpetuating the over-surveillance of birthing people.
Principle 9: Respect and support your data sources.
Much of birthing people’s knowledge about birthing comes from friends, family, and their community (both offline and online). Generations of birthing people have passed down oral stories, written books, and generated huge amounts of online content about pregnancy, labor, and the postpartum period. NLP studies and tools benefit from these datasets, using them as training examples and extracting information from them. But these NLP studies and tools should be designed to avoid supplanting systems of support that are already thriving, and the ownership of this data should be considered.
Beware the “paradox of reuse” in which the creation of an automated tool removes incentives for people to continue creating the training data that powers that tool. Give credit to the data sources. Use proper attribution, as this not only respects people’s work but also builds trust from your users and supports auditing of your system. Where possible, provide back links that lead your users to re-engage with the data source (e.g., by sharing their own experience in an online birthing community).
Table 4: Guiding principles about who and what is valued.
Appendix C Full survey questions
C.1 Consent
Thank you for participating in this study about the use of AI for maternal healthcare. This study is being led by Dr. Maria Antoniak, Allen Institute for AI, and Dr. Carla Alvarado, AAMC Center for Health Justice. The purpose of this research is to gather perceptions about AI tools from a diverse group of people. This information will be used to create a set of guiding principles for the use of AI for maternal health. We deeply value your lived experience, and we want to include as many voices as possible in the development of these guiding principles.
Participation
Your participation in this study is voluntary and anonymous, and you may refuse to participate before the exercise begins, or discontinue at any time with no penalty.
In this study, you will complete a survey describing your experiences with maternal healthcare and your perceptions of AI tools.
Risks & Benefits
This study asks about your experiences in maternal health. Any questions related to this experience may make you uncomfortable or bring up feelings of stress. You are always free to decline to answer any question or stop your participation at any time.
You will be paid for your participation in this study according to the rates specific on Prolific. We hope to learn about the various perspectives in maternal health and represent these perspectives in the guidelines we develop, and we hope that these guidelines will shape future AI research for maternal health in ways that are useful, ethical, and appropriate for the community.
Privacy
We do not collect names or other personally identifiable information. We do collect your Prolific ID, but we do not have any way to link this with your identity. We will not release your Prolific ID to anyone else.
De-identified data from this survey may be shared with the research community at large to advance science and health. We will remove or code any personal information that could identify you before files are shared with other researchers to ensure that, by current scientific standards and known methods, no one will be able to identify you from the information we share. Responses will be stored securely by AAMC and Allen Institute for AI with limited access controls to limit exposure to those with a need to know.
Questions
If you have questions about this study, please reach out to Dr. Carla Alvarado (calvarado@aamc.org) or Dr. Maria Antoniak (mariaa@allenai.org), or send us a message us on Prolific.
C.2 Screening Questions
  1. 1.
    What is your Prolific ID?
  2. 2.
    Do you consent to participate in this study?
  3. 3.
    [Birthing People Only] Have you given birth in the last five years (2018-2023)?
  4. 4.
    [Healthcare Workers Only] Have you ever worked in a healthcare profession?
C.3 Trust
  1. 1.
    Generally speaking, would you say that most people can be trusted, or that you need to be very careful in dealing with people? [Likert scale from (1) need to be very careful to (5) most can be trusted]
  2. 2.
    Generally speaking, would you say that most healthcare providers can be trusted, or that you need to be very careful in dealing with healthcare providers? [Likert scale from (1) need to be very careful to (5) most can be trusted]
C.4 Familiarity and perceptions of AI
In the following sections, we’ll be asking about your familiarity and perceptions of AI systems. Some examples of AI systems include:
  • A chatbot used by people to answer general questions
  • An AI system used by a medical clinic to predict whether a patient has a disease
  • An AI system used by a bank to predict whether an applicant will repay a loan
  • An AI system used by a marketing company to match ads to viewers
  • An AI system used by a streaming company to recommend movies to users
  1. 1.
    How familiar are you with NLP, machine learning, and/or AI? [Not at all familiar, Heard of it but never sought more information, Read about it, Occasionally used tools relying on these methods, Extensively used tools relying on these methods, Trained or finetuned my own models]
  2. 2.
    Have you ever used an AI chatbot like ChatGPT? [Likert scale from (1) never to (5) all the time]
  3. 3.
    AI can eliminate a lot of tedious work for people. [Likert scale from (1) strongly disagree to (5) strongly agree]
  4. 4.
    The overuse of AI may be harmful and damaging to humans. [Likert scale from (1) strongly disagree to (5) strongly agree]
  5. 5.
    Life will be easier and faster with AI. [Likert scale from (1) strongly disagree to (5) strongly agree]
  6. 6.
    AI turns people into just another number. [Likert scale from (1) strongly disagree to (5) strongly agree]
C.5 Values
These definitions are drawn directly from Jakesch et al. [20].
Fairness
A fair NLP system treats all people equally. Developers of fair NLP systems ensure, as far as possible, that the system does not reinforce biases or stereotypes. A fair system works equally well for everyone independent of their race, gender, sexual orientation, and ability.
Privacy
An NLP system that respects people’s privacy implements strong privacy safeguards. Developers of privacy-preserving NLP systems minimize, as far as possible, the collection of sensitive data and ensure that the NLP system provides notice and asks for consent.
Sustainability
A sustainable NLP system preserves the environmental quality of current and future generations. Developers of sustainable NLP systems minimize, as far as possible, electricity use and reduce waste.
Inclusiveness
Inclusive NLP systems empower everyone and engage all people. Developers of inclusive AI systems consider, as far as possible, the needs of people who might otherwise be excluded or marginalized.
Safety
A safe NLP system performs reliably and safely. Developers of safe NLP systems implement strong safety measures. They anticipate and mitigate, as far as possible, physical, emotional, and psychological harms that the system might cause.
Social good
An NLP system that promotes social good supports, as far as possible, human well-being and flourishing, peace and happiness, and the creation of socio-economic opportunities.
Dignity
An NLP system that respects human dignity upholds the inherent worth of every individual. In addition to respective legislation, developers ensure that the system respects human rights and does not diminish human dignity.
Performance
A high-performing NLP system consistently produces good predictions, inferences or answers. Developers of high-performing NLP systems ensure, as far as possible, that the system’s results are useful, accurate and produced with minimal delay.
Accountability
An accountable NLP system has clear attributions of responsibilities and liability. Developers and operators of accountable AI systems are, as far as possible, held responsible for their impacts. An accountable system also implements mechanisms for appeal and recourse.
Transparency
A transparent NLP system produces decisions that people can understand. Developers of transparent AI systems ensure, as far as possible, that users can get insight into why and how a system made a decision or inference.
Human autonomy
An NLP system that respects people’s autonomy avoids reducing their agency. Developers of autonomy-preserving AI systems ensure, as far as possible, that the system provides choices to people and preserves or increases their control over their lives.
Solidarity
A solidary NLP system does not increase inequality and leaves no one behind. Developers of solidary AI systems ensure, as far as possible, that the prosperity as well as the burdens created by NLP are shared by all.
  1. 1.
    Please select any five values from the list below that you think are most important for NLP systems for maternal health. [Fairness, Privacy, Sustainability, Inclusiveness, Safety, Social good, Dignity, Performance, Accountability, Transparency, Human autonomy, Solidarity]
  2. 2.
    (Optional) Was anything missing from the list of values above? If so, describe here.
C.6 Information Seeking Behavior
  1. 1.
    When I have questions about maternal health, I often turn to this resource to assist me.
    [Likert scale from (1) strongly disagree to (5) strongly agree for each of the following options]
    • Peer-reviewed scientific publications (for example: Journal of the American Medical Association)
    • Online communities and forums (for example: Facebook groups, Reddit groups)
    • Social media (for example: Instagram, TikTok, YouTube)
    • Government resources (for example: FDA, CDC websites and announcements)
    • News articles (for example: CNN, WSJ)
    • Textbooks, books
    • Expert-written websites (for example: WebMD, Mayo Clinic)
    • Curated medical resources (for example: UpToDate)
  2. 2.
    (Optional) Outside of the resources listed above, are there are other resources you turn to for help in answering questions about maternal healthcare?
  3. 3.
    When I have questions about maternal health, I often turn to this person to assist me.
    [Likert scale from (1) strongly disagree to (5) strongly agree for each of the following options]
    • OB/GYN
    • Nurse
    • Midwife
    • Doula
    • Community health worker
    • Healthcare reseracher
    • Government workers
    • Friends & family
    • Colleagues
  4. 4.
    (Optional) Outside of the people listed above, are there are other people you turn to for help in answering questions about maternal healthcare?
C.7 Effects of NLP/AI
  1. 1.
    This person would benefit from an NLP/AI chatbot in their work.
    [Likert scale from (1) strongly disagree to (5) strongly agree for each of the following options]
    • OB/GYN
    • Nurse
    • Midwife
    • Doula
    • Community health worker
    • Healthcare reseracher
    • Government workers
    • Friends & family
    • Colleagues
  2. 2.
    [Workshop Only] What impact do you expect an NLP chatbot would have on your work? [Likert scale from (1) mostly harms to (5) mostly benefits]
  3. 3.
    [Birthing People Only] I would want my maternal healthcare provider to use AI systems. [Likert scale from (1) strongly disagree to (5) strongly agree]
  4. 4.
    [Birthing People Only] I would want my maternal healthcare provider to tell me if they use AI systems. [Likert scale from (1) strongly disagree to (5) strongly agree]
  5. 5.
    [Birthing People Only] I would want my maternal healthcare provider to tell me if they use AI systems. [Likert scale from (1) strongly disagree to (5) strongly agree]
  6. 6.
    [Birthing People Only] How often do/did you consult internet resources for pregnancy-related questions during your pregnancy? [Likert scale from (1) never to (5) frequently]
  7. 7.
    [Birthing People Only] If so, what were these pregnancy-related questions about? (check all that apply) [medical worries/concerns during pregnancy, mental health worries/concerns during pregnancy, the pregnancy process in general, labor and delivery, birth control option after delivery, the postpartum process, taking care of a newborn, medical worries/concerns after pregnancy, mental health worries/concerns after pregnancy, Other (write in)]
C.8 Demographics
All of these questions were optional.
  1. 1.
    Which gender(s) do you identify with? [woman, man, non-binary, prefer not to disclose, other (write in)]
  2. 2.
    What is your age range? [$<$18, 18-24, 25-34, 35-44, 45-54, 55-64, 65-74, 74+]
  3. 3.
    Describe your race/ethnicity (check all that apply) [African-American/Black, Middle Eastern/North African, Native American/Alaska Native/First Nations, Pacific Islander, Hispanic/Latino/a/x, East Asian (including Chinese, Japanese, Korean, Mongolian, Tibetan, and Taiwanese), South Asian (including Bangladeshi, Bhutanese, Indian, Nepali, Pakistani, and Sri Lankan), Southeast Asian (including Burmese, Cambodian, Filipino, Hmong, Indonesian, Laotian, Malaysian, Mien, Singaporean, Thai, and Vietnamese), White, Other (write in)]
  4. 4.
    Where do you live in the U.S.? [Northeast, Midwest, South, West, I do not live in the U.S.]
  5. 5.
    [Workshop Only] What is your education level? [Trade school (for example: AA, AS), College (for example: BA, BS), Master’s degree (for example: MS, MA, MPH), Medical degree (for example: MD, DO), Other professional degree (for example: JD), PhD]
  6. 6.
    [Workshop Only] Describe your professional background. Check every profession that has ever applied to you. [Community health worker/Promotoras, Community-based organization (non-profit), Doula, Certified Midwife, Certified Nurse Midwife, L&D Nurse, OB/GYN, Other perinatal health care provider, Other perinatal support (non-health care) provider, Non-perinatal health care provider, Health care management/administration, Health care services researcher, Population Health/Public Health researcher, Local/State public health entity, Federal Government employee in public health (for example: CDC, HHS, CMS and other related units)], AI / machine learning / NLP researcher or engineer, Other (write in)
  7. 7.
    [Workshop Only] Have you ever sought maternal or reproductive support from a healthcare provider (for example: pregnancy support, contraception support)? [yes, no, unsure, not applicable]
  8. 8.
    [Prolific Only] Do you or have you ever worked in healthcare (in a hospital or clinic, as a community health worker, as a public health researcher, or in any other capacity)? [yes, no]
  9. 9.
    [Prolific Only] Do you or have you ever worked in maternal/perinatal healthcare (as a nurse, midwife, researcher, or in any other capacity)? [yes, no]
  10. 10.
    Feel free to share any additional comments or feedback about this survey here. If anything was confusing or unclear, we’d love to know so that we can improve this survey in the future.